# Guided Discrete Diffusion for Electronic Health Record Generation

### v1

@@ Abstract
Electronic health records (EHRs) are a pivotal data source that enables numerous applications in computational medicine, e.g., disease progression prediction, clinical trial design, and health economics and outcomes research. Despite wide usability, their sensitive nature raises privacy and confidentially concerns, which limit potential use cases. To tackle these challenges, we explore the use of generative models to synthesize artificial, yet realistic EHRs. While diffusion-based methods have recently demonstrated state-of-the-art performance in generating other data modalities and overcome the training instability and mode collapse issues that plague previous GAN-based approaches, their applications in EHR generation remain underexplored. The discrete nature of tabular medical code data in EHRs poses challenges for high-quality data generation, especially for continuous diffusion models. To this end, we introduce a novel tabular EHR generation method, EHR-D3PM, which enables both unconditional and conditional generation using the discrete diffusion model. Our experiments demonstrate that EHR-D3PM significantly outperforms existing generative baselines on comprehensive fidelity and utility metrics while maintaining less membership vulnerability risks. Furthermore, we show EHR-D3PM is effective as a data augmentation method and enhances performance on downstream tasks when combined with real data.
@@ 1 Introduction
Electronic health records (EHRs) are a rich and comprehensive data source, enabling numerous applications in computational medicine including the development of models for disease progression prediction and clinical event medical models (Li et al., 2020; Rajkomar et al., 2018), clinical trial design (Bartlett et al., 2019) and health economics and outcome research (Padula et al., 2022). In particular, many existing disease prediction models primarily utilize tabular formats, often transforming longitudinal EHR data into binary or categorical forms, rather than employing time-series forecasting methods (Lee et al., 2022; Huang et al., 2021; Rao et al., 2023; Debal and Sitote, 2022). However, the sensitive nature of EHRs, which includes confidential medical data, poses challenges for their broad use due to privacy concerns and patient confidentiality requirements (Hodge Jr et al., 1999). In addition to these concerns, data scarcity also restricts their potential use in applications for rare medical conditions. To address these challenges, we consider using generative models to synthesize artificial, but realistic EHRs, which has recently emerged as a crucial research area for advancing applications of machine learning to healthcare and other industries with privacy and data scarcity challenges.
@@ 2 Related Work
EHR Synthesis. Various methods have been developed for generating synthetic EHR data. Buczak et al. (2010) proposed an early data-driven approach for creating synthetic EHRs, but their approach offers limited flexibility has privacy concerns. Recently, GANs have become prominent in EHR generation, including medGAN Choi et al. (2017b), medBGAN (Baowaly et al., 2018), EHRWGAN (Zhang et al., 2019), and CorGAN Torfi and Fox (2020b). GAN-based methods offer significant improvement in the quality of synthetic EHRs, but often face issues related to training instability and mode collapse (Thanh-Tung et al., 2018), restricting their wide use and the diversity of generated data. To address this, other methods, including variational auto-encoders (Biswal et al., 2020) and language models (Wang and Sun, 2022), have been explored. Very recently, MedDiff (He et al., 2023) and EHRdiff (Yuan et al., 2023) considered using diffusion models and proposed sampling techniques for high-quality EHR generation. Ceritli et al. (2023) further extended the diffusion model to mixed-type EHRs. In this paper, we focus on developing a guided discrete diffusion model specifically designed for generating tabular medical codes in EHRs and improving the generation of codes for rare conditions, which previous methods struggle at.
Discrete Diffusion Models. The study of discrete diffusion models was pioneered by Sohl-Dickstein et al. (2015), which explored diffusion processes in binary random variables. The approach was further developed by Ho et al. (2020); Song et al. (2020), which incorporates categorical random variables using transition matrices with uniform probabilities. Subsequently, Austin et al. (2021) introduced a generalized framework named Discrete Denoising Diffusion Probabilistic Models (D3PMs) for categorical random variables, effectively combining discrete diffusion models with Masked Language Models (MLMs). Recent advancements in this field include the introduction of editing-based operations (Jolicoeur-Martineau et al., 2021; Reid et al., 2022), auto-regressive diffusion models (Hoogeboom et al., 2021a; Ye et al., 2023), a continuous-time structure (Campbell et al., 2022), strides in generation acceleration (Chen et al., 2023), and the application of neural network analogs for learning purposes (Sun et al., 2022). In this paper, we focus on the D3PMs with multinomial distribution.
@@ 3 Background
$=\mathbb{E}_{q(\mathbf{x}_{1}|\mathbf{x}_{0})}[\log p_{\bm{\theta}}(\mathbf{x}_{0}|\mathbf{x}_{1})]$
$-\sum_{t=2}^{T}\mathbb{E}_{q(\mathbf{x}_{t}|\mathbf{x}_{0})}[\mathrm{KL}(q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})\|p_{\bm{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t}))$
$-\mathbb{E}_{q(\mathbf{x}_{T}|\mathbf{x}_{0})}\mathrm{KL}(q(\mathbf{x}_{T}|\mathbf{x}_{0})\|p_{\bm{\theta}}(\mathbf{x}_{T})),$
@@
where $\beta_{t}$ is the variance schedule determined by a pre-specified corruption schedule. The Gaussian diffusion process has achieved great success in continuous-valued applications like image generation (Ho et al., 2020; Song et al., 2020). Recently, it has been applied to tabular EHR data generation (He et al., 2023; Yuan et al., 2023).
@@ 4.1 Problem Formulation
We consider medical coding data in EHRs, such as ICD codes, which are standardized codes published by the World Health Organization that correspond to specific medical diagnoses and procedures (Slee, 1978). While we focus on ICD codes specifically, our approach can be used with other medical coding data, e.g., CPT, NDC and LOINC codes. For a given (usually high dimensional) set $\Omega$ of ICD codes of interest, we encode the set as $N:=|\Omega|$ categories $\{1,2,\ldots,N\}$. A sample patient EHR $\mathbf{x}$ is then encoded as a sequence of $N$ tokens $\mathbf{x}=[\mathbf{x}^{(1)},\ldots,\mathbf{x}^{(N)}]$, where each token $\mathbf{x}^{(i)}\in\{0,1\}^{2}$ is a one-hot function. $\mathbf{x}^{(i)}$ represents the occurrence of the $i$-th ICD code in the patient EHR. In particular, $\mathbf{x}^{(i)}=[1,0]$ represents occurrence of code and $\mathbf{x}^{(i)}=[0,1]$ represents its absence. We assume a sufficiently large set of patient EHRs is available to train a multinomial diffusion model to generate artificial encoded patient EHRs sequences $\mathbf{x}^{\prime}$.
@@ 4.2 Unconditional Generation
The reverse sampling procedure uses the predictor $p_{\bm{\theta}}(\cdot|\mathbf{x}_{t})$ with the following neural network architecture:
$\mathbf{z}_{0,t}$ $=\mathbf{x}_{t}=[\mathbf{x}_{t}^{(1)},\ldots,\mathbf{x}_{t}^{(N)}]$
$\mathbf{z}_{l,t}^{\prime\prime}$ $=[\mathbf{z}_{l-1,t}^{(1)},\ldots,\mathbf{z}_{l-1,t}^{(N)}]+\mathbf{E}_{\mathrm{pos}}+\mathbf{E}_{\mathrm{time}},$
$\mathbf{z}_{l,t}^{\prime}$ $=\mathrm{LinMSA}(\mathrm{LN}(\mathbf{z}^{\prime\prime}_{l-1,t}))+\mathbf{z}^{\prime\prime}_{l-1,t},$
$\mathbf{z}_{L,t}$ $=\mathrm{ParallelMLP}(\mathrm{LN}(\mathbf{z}_{L-1,t}^{\prime})),$
$\mathrm{Output}$ $=[\mathrm{softmax}(\mathbf{z}_{L,t}^{(1)}),\ldots,\mathrm{softmax}(\mathbf{z}_{L,t}^{(N)})]$ (4)
where $\mathbf{E}_{\mathrm{pos}},\mathbf{E}_{\mathrm{time}}\in\mathbb{R}^{2\times D}$ represents the position embedding and time embedding respectively, the variable $l$ indexes the layers belonging to the set $\{1,\ldots,L\}$, $\mathrm{LinMSA}$ refers to the linear-time multi-head self-attention block proposed by Wang et al. (2020) and $\mathrm{LN}$ is an abbreviation for layer normalization. For each dimension of $\widehat{\mathbf{x}}_{0}$, we apply a multilayer perceptron layer to obtain the logit, abbreviated as ParallelMLP. The $\mathrm{softmax}$ function transforms the last-layer latent variable $\mathbf{z}_{L}^{i}$ into the conditional probability $p_{\bm{\theta}}(\cdot|\mathbf{x}_{t})$, serving as the final softmax layer. The details of our denoise model are provided in Fig. 6 in Appendix A.2.
@@ 4.3 Conditional Generation with Classifier Guidance
The goal of conditional generation is to generate $p_{\bm{\theta}}(\mathbf{x}|\mathbf{c})$ close to $q_{\mathrm{data}}(\mathbf{x}|\mathbf{c})$, where $\mathbf{c}$ denotes a context, such as the presence of a single or group of ICD codes in a patient EHR. $\mathbf{c}$ is not available at training time, but we assume access to a classifier $p(\mathbf{c}|\mathbf{x})$ that is close to the conditional distribution $q_{\mathrm{data}}(\mathbf{c}|\mathbf{x})$. Then given an unconditional EHR generator $p_{\bm{\theta}}(\mathbf{x})$ and classifier $p(\mathbf{c}|\mathbf{x})$, we propose a training-free conditional generator as follows
@@
To sample from (5), we apply the following guided multinomial diffusion procedure:
@@
where $\widehat{\mathbf{x}}_{0}$ is the latent variable that predicts $\mathbf{x}_{0}$. Since $\widehat{\mathbf{x}}_{0}$ lies in a discrete space, we cannot directly use the Langevin dynamics in the space of $\widehat{\mathbf{x}}_{0}$. However, the last-layer latent variable $\mathbf{z}_{L,t}$ in (4) (before the softmax layer) lies in a continuous space. We have the following:
@@
Therefore, we can use the plug-and-play method (Dathathri et al., 2019) with the latent space $\mathbf{z}_{L}$, which has been recently employed in text generation (Dathathri et al., 2019) and protein design (Gruver et al., 2023). In particular, we introduce a modified latent variable $\mathbf{y}^{(k)}$ for $\mathbf{z}_{L,t}$, which is initialized as $\mathbf{y}^{(0)}\leftarrow\mathbf{z}_{L,t}$. Then we iterative apply the following update with Langevin dynamics
@@
where the energy function $V_{\bm{\theta}}(\mathbf{y}^{(k)})=\log(p(\mathbf{c}|\mathbf{y}^{(k)}))=\log\big{(}\sum_{\widehat{\mathbf{x}}_{0}}p_{\bm{\theta}}(\widehat{\mathbf{x}}_{0}|\mathbf{y}^{(k)})p(\mathbf{c}|\widehat{\mathbf{x}}_{0})\big{)}$ and $\mathcal{D}_{\mathrm{KL}}(\mathbf{y}^{(k)})=\lambda\mathrm{KL}(p_{\theta}(\widehat{\mathbf{x}}_{0}|\mathbf{y}^{(k)})||p_{\theta}(\widehat{\mathbf{x}}_{0}|\mathbf{y}^{(0)}))$ is the Kullback–Leibler ($\mathrm{KL}$) divergence for regularization of the guided Markov transition. The gradient of the energy term $\nabla_{\mathbf{y}^{(k)}}V_{\bm{\theta}}$ drives the hidden state $\mathbf{y}^{(k)}$ towards high probability of $p(\mathbf{c}|\mathbf{y}^{(k)})$. The gradient of the regularization term $\nabla_{\mathbf{y}^{(k)}}\mathcal{D}_{\mathrm{KL}}$ ensures the guided transition distribution still maximizes the likelihood of the diffusion model. For a more detailed discussion, see Appendix C.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) Spearm corr = 0.82 (b) Spearm corr = 0.83 (c) Spearm corr = 0.75 (d) Spearm corr = 0.94
Refer to caption Refer to caption Refer to caption Refer to caption
(e) Spearm corr = 0.79 (f) Spearm corr = 0.80 (g) Spearm corr = 0.71 (h) Spearm corr = 0.93
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Density comparison of per-record feature number for synthetic and real data for the MIMIC dataset. The number of features per record is the sum of ICD codes present in each sample. The number of bins is 40, and the range of feature number values is (0, 40).
Refer to caption Refer to caption Refer to caption Refer to caption
(a) Spearm corr = 0.91 (b) Spearm corr = 0.98 (c) Spearm corr = 0.85 (d) Spearm corr = 0.99
Refer to caption Refer to caption Refer to caption Refer to caption
(e) Spearm corr = 0.87 (f) Spearm corr = 0.91 (g) Spearm corr = 0.79 (h) Spearm corr = 0.97

### v2

@@ Abstract
Electronic health records (EHRs) are a pivotal data source that enables numerous applications in computational medicine, e.g., disease progression prediction, clinical trial design, and health economics and outcomes research. Despite wide usability, their sensitive nature raises privacy and confidentially concerns, which limit potential use cases. To tackle these challenges, we explore the use of generative models to synthesize artificial, yet realistic EHRs. While diffusion-based methods have recently demonstrated state-of-the-art performance in generating other data modalities and overcome the training instability and mode collapse issues that plague previous GAN-based approaches, their applications in EHR generation remain underexplored. The discrete nature of tabular medical code data in EHRs poses challenges for high-quality data generation, especially for continuous diffusion models. To this end, we introduce a novel tabular EHR generation method, EHR-D3PM, which enables both unconditional and conditional generation using the discrete diffusion model. Our experiments demonstrate that EHR-D3PM significantly outperforms existing generative baselines on comprehensive fidelity and utility metrics while maintaining less attribute and membership vulnerability risks. Furthermore, we show EHR-D3PM is effective as a data augmentation method and enhances performance on downstream tasks when combined with real data.
@@ 1 Introduction
Electronic health records (EHRs) are a rich and comprehensive data source, enabling numerous applications in computational medicine including the development of models for disease progression prediction and clinical event medical models (Li et al., 2020; Rajkomar et al., 2018), clinical trial design (Bartlett et al., 2019) and health economics and outcome research (Padula et al., 2022). In particular, many existing disease prediction models primarily utilize tabular formats, often transforming longitudinal EHR data into binary or categorical forms, rather than employing time-series forecasting methods (Lee et al., 2022a; Huang et al., 2021; Rao et al., 2023; Debal and Sitote, 2022a). However, the sensitive nature of EHRs, which includes confidential medical data, poses challenges for their broad use due to privacy concerns and patient confidentiality requirements (Hodge Jr et al., 1999). In addition to these concerns, data scarcity also restricts their potential use in applications for rare medical conditions. To address these challenges, we consider using generative models to synthesize artificial, but realistic EHRs, which has recently emerged as a crucial research area for advancing applications of machine learning to healthcare and other industries with privacy and data scarcity challenges.
@@ 2 Related Work
EHR Synthesis. Various methods have been developed for generating synthetic EHR data. Buczak et al. (2010) proposed an early data-driven approach for creating synthetic EHRs, but their approach offers limited flexibility has privacy concerns. Recently, GANs have become prominent in EHR generation, including medGAN Choi et al. (2017b), medBGAN (Baowaly et al., 2018), EHRWGAN (Zhang et al., 2019), and CorGAN Torfi and Fox (2020b). GAN-based methods offer significant improvement in the quality of synthetic EHRs, but often face issues related to training instability and mode collapse (Thanh-Tung et al., 2018), restricting their wide use and the diversity of generated data. To address this, other methods, including variational auto-encoders (Biswal et al., 2020) and language models (Wang and Sun, 2022), have been explored. Very recently, MedDiff (He et al., 2023b) considered using diffusion models and proposed sampling techniques for high-quality EHR generation. Ceritli et al. (2023); Yuan et al. (2023) further extended the diffusion model to mixed-type EHRs. In this paper, we focus on developing a guided discrete diffusion model tailored specifically for generating tabular medical codes in Electronic Health Records (EHRs), which has wide-ranging applications in healthcare (Debal and Sitote, 2022b; Lee et al., 2022b). Our model aims to improve the generation of ICD codes for rare conditions, which has been a challenge for previous methods.
Discrete Diffusion Models. The study of discrete diffusion models was pioneered by Sohl-Dickstein et al. (2015), which explored diffusion processes in binary random variables. The approach was further developed by Ho et al. (2020); Song et al. (2020), which incorporates categorical random variables using transition matrices with uniform probabilities. Subsequently, Austin et al. (2021) introduced a generalized framework named Discrete Denoising Diffusion Probabilistic Models (D3PMs) for categorical random variables, effectively combining discrete diffusion models with Masked Language Models (MLMs). Recent advancements in this field include the introduction of editing-based operations (Jolicoeur-Martineau et al., 2021; Reid et al., 2022), auto-regressive diffusion models (Hoogeboom et al., 2021a; Ye et al., 2023), a continuous-time structure (Campbell et al., 2022), strides in generation acceleration (Chen et al., 2023), and the application of neural network analogs for learning purposes (Sun et al., 2022). Diffusion models (Kim et al., 2022; Lee et al., 2023; Kotelnikov et al., 2023; Zhang et al., 2023) are recently developed for tabular data generation but not specifically designed for sparse and high dimensional EHR generation. In this paper, we are based on D3PMs with multinomial distribution for EHR generation.
@@ 3 Background
$=\mathbb{E}_{q(\mathbf{x}_{1}|\mathbf{x}_{0})}[\log p_{\bm{\theta}}(\mathbf{x}_{0}|\mathbf{x}_{1})]-\sum_{t=2}^{T}\mathbb{E}_{q(\mathbf{x}_{t}|\mathbf{x}_{0})}[\mathrm{KL}(q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})\|p_{\bm{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t}))$
$\qquad-\mathbb{E}_{q(\mathbf{x}_{T}|\mathbf{x}_{0})}\mathrm{KL}(q(\mathbf{x}_{T}|\mathbf{x}_{0})\|p_{\bm{\theta}}(\mathbf{x}_{T})),$
@@
where $\beta_{t}$ is the variance schedule determined by a pre-specified corruption schedule. The Gaussian diffusion process has achieved great success in continuous-valued applications like image generation (Ho et al., 2020; Song et al., 2020). Recently, it has been applied to tabular EHR data generation (He et al., 2023b; Yuan et al., 2023).
@@ 4.1 Problem Formulation
We consider medical coding data in EHRs, such as diagnose codes (ICD), procedure codes (CPT) and medication codes (GEN), which are standardized codes published by the World Health Organization that correspond to specific medical diagnoses and procedures (Slee, 1978). In a wide application of medical domains, it is common to convert continuous variables into discrete ones to enhance the performance of prediction models (Rasmy et al., 2021; Hill et al., 2023). For example, Hill et al. (2023) converts continuous lab codes to discrete tokens based on deciles of each LOINC code; prediction models built on learnt representation of tokenized discrete data significantly outperform ML models which are directly trained on the mixed-type tabular data. Therefore, our paper focus on the generation of discrete medical codes, e.g., ICD, CPT, GEN and LOINC codes.
For a given (usually high dimensional) set $\Omega$ of medical codes of interest, we encode the set as $N:=|\Omega|$ categories $\{1,2,\ldots,N\}$. A sample patient EHR $\mathbf{x}$ is then encoded as a sequence of $N$ tokens $\mathbf{x}=[\mathbf{x}^{(1)},\ldots,\mathbf{x}^{(N)}]$, where each token $\mathbf{x}^{(i)}\in\{0,1\}^{2}$ is a one-hot function. $\mathbf{x}^{(i)}$ represents the occurrence of the $i$-th medical code in the patient EHR. In particular, $\mathbf{x}^{(i)}=[1,0]$ represents occurrence of code and $\mathbf{x}^{(i)}=[0,1]$ represents its absence. We assume a sufficiently large set of patient EHRs is available to train a multinomial diffusion model to generate artificial encoded patient EHRs sequences $\mathbf{x}^{\prime}$.
@@ 4.2 Unconditional Generation
The reverse sampling procedure uses the predictor $p_{\bm{\theta}}(\cdot|\mathbf{x}_{t})$ with the following neural network architecture with $\mathbf{z}_{0,t}=\mathbf{x}_{t}=[\mathbf{x}_{t}^{(1)},\ldots,\mathbf{x}_{t}^{(N)}]$,
$\mathbf{z}_{l,t}^{\prime}$ $=\mathbf{z}_{l-1,t}+\mathbf{E}_{\mathrm{pos}}+\mathbf{E}_{\mathrm{time}},$ $\mathbf{z}_{l,t}$ $=\mathrm{LinMSA}(\mathrm{LN}(\mathbf{z}^{\prime}_{l-1,t}))+\mathbf{z}^{\prime}_{l-1,t},$
$\mathbf{z}_{L,t}$ $=\mathrm{MLP}(\mathrm{LN}(\mathbf{z}_{L-1,t})),$ $\mathrm{Output}$ $=\mathrm{Softmax}(\mathbf{z}_{L,t}),$ (4)
where $\mathbf{E}_{\mathrm{pos}},\mathbf{E}_{\mathrm{time}}\in\mathbb{R}^{2\times D}$ represents the position embedding and time embedding respectively, the variable $l$ indexes the layers belonging to the set $\{1,\ldots,L\}$. $\mathrm{LinMSA}$ refers to the efficient multi-head self-attention block proposed by Wang et al. (2020), which has linear complexity with respect to the input dimension. $\mathrm{LN}$ is an abbreviation for layer normalization. For each dimension of $\widehat{\mathbf{x}}_{0}$, we apply a multilayer perceptron layer to obtain the logit, abbreviated as ParallelMLP. The $\mathrm{softmax}$ function transforms the last-layer latent variable $\mathbf{z}_{L}^{i}$ into the conditional probability $p_{\bm{\theta}}(\cdot|\mathbf{x}_{t})$, serving as the final softmax layer. The details of our denoise model are provided in Figure 4 in Appendix A.2.
@@ 4.3 Conditional Generation with Classifier Guidance
The goal of conditional generation is to generate $p_{\bm{\theta}}(\mathbf{x}|\mathbf{c})$ close to $q_{\mathrm{data}}(\mathbf{x}|\mathbf{c})$, where $\mathbf{c}$ denotes a context, such as the presence of a single or group of medical codes in a patient’s Electronic Health Record (EHR). Since $\mathbf{c}$ is not available during the training process, we cannot train the generator $p_{\bm{\theta}}(\mathbf{x}|\mathbf{c})$ directly. However, we assume access to a classifier $p(\mathbf{c}|\mathbf{x})$ that approximates the conditional distribution $q_{\mathrm{data}}(\mathbf{c}|\mathbf{x})$. Given an unconditional EHR generator $p_{\bm{\theta}}(\mathbf{x})$ and classifier $p(\mathbf{c}|\mathbf{x})$, we can propose a training-free conditional generator as follows:
@@
To sample from equation (5), the most popular approach is applying Langevin sampling on the unnormalized joint density while injecting Gaussian noise (Ho and Salimans, 2022). However, there is a challenge in multinomial diffusion as the state $\mathbf{x}$ lies in a discrete space. In particular, the multinomial diffusion procedure is as follows:
@@
where $\widehat{\mathbf{x}}_{0}$ is the latent variable that predicts $\mathbf{x}_{0}$. It is not possible to directly apply Langevin dynamics in the space of $\widehat{\mathbf{x}}_{0}$. However, the last-layer latent variable $\mathbf{z}_{L,t}$ in Equation (4) (before the softmax layer) lies in a continuous space. We have the following:
@@
Therefore, we can utilize the plug-and-play method, initially employed in text generation (Dathathri et al., 2019) and recently applied to protein design (Gruver et al., 2023), with the latent space $\mathbf{z}_{L}$. In particular, we introduce a modified latent variable $\mathbf{y}^{(k)}$ for $\mathbf{z}_{L,t}$, which is initialized as $\mathbf{y}^{(0)}\leftarrow\mathbf{z}_{L,t}$. This modification allows us to iteratively update the latent variable using Langevin dynamics, guiding it towards a desired target while maintaining the learned structure of the diffusion model. The iterative update is applied as follows:
@@
where the energy function $V_{\bm{\theta}}(\mathbf{y}^{(k)})=\log(p(\mathbf{c}|\mathbf{y}^{(k)}))=\log\big{(}\sum_{\widehat{\mathbf{x}}_{0}}p_{\bm{\theta}}(\widehat{\mathbf{x}}_{0}|\mathbf{y}^{(k)})p(\mathbf{c}|\widehat{\mathbf{x}}_{0})\big{)}$ and $\mathcal{D}_{\mathrm{KL}}(\mathbf{y}^{(k)})=\lambda\mathrm{KL}(p_{\theta}(\widehat{\mathbf{x}}_{0}|\mathbf{y}^{(k)})||p_{\theta}(\widehat{\mathbf{x}}_{0}|\mathbf{y}^{(0)}))$ is the Kullback–Leibler ($\mathrm{KL}$) divergence for regularization of the guided Markov transition. The gradient of the energy term $\nabla_{\mathbf{y}^{(k)}}V_{\bm{\theta}}$ drives the hidden state $\mathbf{y}^{(k)}$ towards high probability of $p(\mathbf{c}|\mathbf{y}^{(k)})$, ensuring that the generated samples align with the desired target. The gradient of the regularization term $\nabla_{\mathbf{y}^{(k)}}\mathcal{D}_{\mathrm{KL}}$ ensures that the guided transition distribution still maximizes the likelihood of the diffusion model, preserving the learned structure and preventing excessive deviation from the original model. For a more detailed discussion, see Appendix C.